INFRARED PHOTODISSOCIATION SPECTROSCOPY OF M^{+}-(C_{2}H_{2})_{n}) CLUSTERS^{a}

\underline{T. JAEGER}, R. S. WALTERS, and M. A. DUNCAN , University of Georgia, Department of Chemistry, Athens, GA, 30602- 2556.

Metal ion-acetylene complexes of the general form M^{+}-(C_{2}H_{2})_{n} are synthesized in the gas phase by a pulsed nozzle laser ablation cluster source. These ions are then mass selected in a reflectron time-of-flight mass spectrometer where they are excited and undergo multiphoton absorption from a tunable infrared OPO/OPA laser system. Efficient fragmentation occurs by the loss of whole C_{2}H_{2} units and this process is enhanced on vibrational resonances near the symmetric and asymmetric C-H stretch frequencies of acetylene. Infrared photofragmentation spectra are obtained by monitoring the fragment yield versus energy of the tunable infrared laser in the 3000 cm^{−1} to 3400 cm^{−1} region. the observed infrared band positions are red-shifted from those of free acetylene consistent with the weakening of the carbon-carbon bond energy by π electron withdrawal. Interpretation of the spectra for M^{+} = Ni^{+} and Co^{+} will be discussed.

aResearch sponsored by the U.S. Department of Energy